OpenStudy (anonymous):

cos (2pi/9) cos (pi/9) - sin (2pi/9) sin(pi/9)

OpenStudy (aum):

cos(A)cos(B) - sin(A)sin(B) = cos(A+B) Here, A = 2pi/9 and B = pi/9

OpenStudy (cwrw238):

so now you need to simplify 2pi/9 + pi/9 what does that come to?